OpenStudy (anonymous):

x^2+3x-4=0 solve by completing the square what is x=

sam (.sam.):

x^2+3x-4=0 (x+3/2)^2-9/4-4=0 (x+3/2)^2-25/4=0

OpenStudy (anonymous):

sam is correct^_^